solve the following equation

  1. d2y/dx2-2dy/dx+5y=extan2x
1
Expert's answer
2021-07-19T10:56:29-0400

Solution

To get general solution of homogeneous equation y0’’-2y0’+5y0=0 let’s solve characteristic equation r2-2r+5=0.

r1,2=1±2i

So y0(x) = C1exsin(2x)+C2excos(2x), where C1 and C2 are arbitrary constants.

Partial solution x1(t) may be find in the form y1(x) = A(x)ϕ1(x)+ B(x)ϕ2(x), where ϕ1(x)= exsin(2x),  ϕ2(x)= excos(2x) Substituting into equation according to the method of undetermined coefficients  we’ll get

A’ϕ1(x)+ B’ϕ2(x)=0, A’ϕ’1(x)+ B’ϕ’2(x)= extan2x

From this system

A(x)=-∫ϕ2(x) extan2x dx/W(x), B(x)=∫ϕ1(x) extan2x dx/W(x), where W(x) = ϕ1(x) ϕ’2(x) - ϕ2(x) ϕ’1(x)

ϕ’1(x) = exsin(2x)+2 excos(2x), ϕ’2(x) = excos(2x) - 2 exsin(2x)  =>  W(x) = -2e2x  

A(x) = -0.5∫cos(2x)tan(2x)dx = -0.25∫tan(2x)dsin(2x) = 0.25cos(2x)

B(x) = -0.5∫sin(2x)tan(2x)dx = 0.25∫tan(2x)dcos(2x) = 0.25[sin(2x)-ln|(1+sin(2x))/cos(2x)|]

So y1(x) = A(x)ϕ1(x)+ B(x)ϕ2(x) = -0.25 excos(2x) ln|(1+sin(2x))/cos(2x)|]

and solution of DE is

y(x) = y0(x) + y1(x) = C1exsin(2x)+C2excos(2x) - 0.25 excos(2x) ln|(1+sin(2x))/cos(2x)|]